today only, a table is being sold at a 38% discount. The sale price is 217. what was the price yesterday?
Answer by Cromlix(4381)

Hi there,
38% discount = 100% - 38% = 62%
62% = 217
1% = 217/62
100% = 217/62 x 100
100% = 350 (the price yesterday)
Hope this helps :-)